If the marginal costs are constant and zero for a single price monopolist facing the demand curve P = 10 - Q, what will profits be if fixed costs are 12?
Answer #1

P= 10 - Q

TR = PQ = 10Q - Q^2

MR = 10 - 2Q

MR= MC

10-2Q = 0

Q= 5

P= 10-5= 5

Profit = 5*5 - 12= 13